# OFDM-Based Active STAR-RIS-Aided Integrated Sensing and Communication Systems

### v1

@@ I-B Paper outline
For the rest of this paper, Section II introduces the structure of the ASTARS-aided ISAC system and the expressions of radar and communication signals. In Section III, we introduce the optimization problem of our system and the design of transmit beamforming. Section IV presents the optimization of the ASTARS phases and amplitudes. We simplify the problem to enable the use of CVX for carrying out the optimization. In Section V, we introduce the procedures for range and velocity estimations by using our optimized transmit beamforming and ASTARS phases and amplitudes. Section VI includes simulations, and the Conclusion is presented in Section VII.
@@ II-A ASTARS model
We consider a downlink ASTARS-aided ISAC system designed for simultaneous radar sensing and communication as Fig. $1$. The DFRC BS transmits a signal to the ASTARS with $Q$ elements, which then directs the beam towards a target in the sensing area. Upon receiving the beam, the target reflects it back along its path, which is then captured by the ASTARS and redirected to the DFRC BS for processing. The DFRC BS employs uniform linear arrays (ULAs) for both sensing and communication, consisting of $M_{T}$ and $M_{R}$ elements, respectively, and totaling $M$ antennas. These arrays are utilized to communicate with $K$ users, each equipped with a single antenna. Notably, these users are situated on the opposite side of the ASTARS, distinct from the sensing target location. We assume that $\mathcal{K}_{t}=\{1,\ldots,K\}$ users are located on the communication region $(t)$, i.e., behind the surface. We assume that all surface elements serve simultaneously all users in $t$ region and the target in $r$ region. The amplitude and phase parameters for the ASTARS elements are defined as follows: The amplitude for $q$th element $\beta_{q}^{i}$ lies within the range $[0,1]$ and the phase $\phi_{q}^{i}$ within $[0,2\pi)$, where the index $i$ denotes either communication ($t$) or sensing ($r$). The entire phase matrix $\boldsymbol{\Psi}_{i}$ is constructed as a diagonal matrix with elements $\alpha_{q}\beta_{q}^{i}e^{j\phi_{q}^{i}}$, resulting in $\boldsymbol{\Psi}_{i}=\mathrm{diag}(\alpha_{1}\beta_{1}^{i}e^{j\phi_{1}^{i}},\ldots,\alpha_{Q}\beta_{Q}^{i}e^{j\phi_{Q}^{i}})\in\mathbb{C}^{Q\times Q}$. The vector $\boldsymbol{\phi}_{i}$ contains the phase elements of each ASTARS element, expressed as $\boldsymbol{\phi}_{i}=[\alpha_{1}\beta_{1}^{i}e^{j\phi_{1}^{i}},\ldots,\alpha_{Q}\beta_{Q}^{i}e^{j\phi_{Q}^{i}}]^{T}$. It is subject to the conditions that for each element $q$, the sum of the squares of the amplitudes for communication and sensing must be equal to one, i.e., ($(\beta_{q}^{t})^{2}+(\beta_{q}^{r})^{2}=1$), the active amplifying coefficient (AAC) $\alpha_{q}$ should not exceed a maximum value $\alpha_{max}$, and the phase $\phi_{q}^{i}$ must fall within the range $[0,2\pi)$ for all $q$. To acquire the OFDM symbol, the transmitted block signals undergo an inverse discrete Fourier transform (IDFT). Following the IDFT, the output is cyclically shifted to mitigate intersymbol interference (ISI) and to incorporate a guard interval. Once the signals traverse multipath channels, it becomes essential to eliminate the guard interval and execute the discrete Fourier transform (DFT) [28], [31]. The assumption here is that the cyclic prefix (CP) length is sufficiently large to disregard any timing and frequency discrepancies.
@@ II-C Signal at the radar receiver
For the channels in the radar sensing area, we consider that the channel between DFRC BS and the ASTARS is $\mathbf{G}\in\mathbb{C}^{Q\times M}$, which is based on Rayleigh fading. The channel between the ASTARS and the target is $\mathbf{h}_{a,T}\in\mathbb{C}^{Q\times 1}$. To simpify the complexity, the direct link from the DFRC BS to the target is blocked in our case like [21]. Hence, the collected echo signal at the DFRC BS from the target can be written as
$\mathbf{y}_{r}(n)=\mathbf{G}^{H}(n)[\boldsymbol{\Psi}_{r}\mathbf{h}_{a,T}(n)\mathbf{h}_{a,T}^{H}(n)\boldsymbol{\Psi}_{r}\mathbf{G}(\mathbf{x}(n)+\mathbf{v}_{1}(n))$ (3)
@@ II-D Communication signal at the users
$=(\mathbf{t}_{k}^{H}(n)\boldsymbol{\Psi}_{t}^{H}(n)\mathbf{G}^{H}(n)+\mathbf{h}_{d,k}^{H}(n))\mathbf{x}(n)+\mathbf{t}_{k}^{H}(n)\boldsymbol{\Psi}_{t}^{H}\mathbf{v}_{1}(n)$
@@
$\mathrm{SINR}_{k}(n)=\frac{|\mathbf{u}_{k}^{H}(n)\mathbf{w}_{c,k}|_{2}^{2}}{|\mathbf{u}_{k}^{H}(n)\mathbf{W}_{r}|_{2}^{2}+\sum_{i\neq k}\mathbf{u}_{k}^{H}(n)\mathbf{R}_{i}\mathbf{u}_{k}+\sigma_{v}^{2}\mathbf{t}_{k}^{H}(n)\boldsymbol{\Psi}_{t}\boldsymbol{\Psi}_{t}^{H}(n)\mathbf{t}_{k}(n)+\sigma^{2}}$ (8)
where $\mathbf{h}_{d,k}(n)\in\mathbb{C}^{M\times 1}$ is the direct link from DFRC BS to users in the $n$th subcarrier; $\mathbf{t}_{k}(n)\in\mathbb{C}^{Q\times 1}$ is the channel between the STAR-RIS and user $k$. $n_{k}(n)\sim\mathcal{CN}(0,\sigma^{2})$ is the AWGN. The SINR for the $k$th user can be given as (8), where $\mathbf{u}_{k}(n)=\mathbf{h}_{d,k}(n)+\mathbf{G}(n)\boldsymbol{\Psi}_{t}\mathbf{t}_{k}(n)$ is the overall channel between users and the DFRC BS. Also, channels $\mathbf{h}_{d,k}$, $\mathbf{G}$, and $\mathbf{t}_{k}$ are based on Rayleigh fading, with large-scale fading coefficients $\widetilde{\beta}_{d,k}$, $\widetilde{\beta}_{G}$, and $\eta_{k}$.
@@ III-A Problem formulation
$\quad s.t.\quad\quad\quad\sum_{k}\mathrm{SINR}_{k}(n)\geq\xi,\forall k$ (9a)
$\quad\quad\quad\quad\quad\quad\mathrm{tr}(\mathbf{W})\leq P_{BS}$ (9b)
@@ III-B Transmit beamforming $\mathbf{W}$ optimization
$\;\quad\mathrm{tr}(\mathbf{W})\leq P_{BS},$
@@
IV Phase coefficients optimization
@@
V Range and velocity estimation
@@
V-A Estimations for range and velocity
@@
V-B MSE for range and velocity estimations
@@
VI Simulations
@@
In this section, we provide simulation results for our proposed ASTARS-aided ISAC systems with OFDM. The DFRC BS is equipped with $M=32$ antennas, which serves $K=4$ users located on the communication area. The ASTARS comprises an equal number of elements in both dimensions, for example, forming a $10\times 10$ array when $Q=100$ elements. In our system, we assume a carrier frequency $3$ GHz with wavelength $\lambda=0.1$ m. We assume the large-scale fading coefficients $\widetilde{\beta}_{d,k}=C_{1}d_{1}^{-\mu_{1}}$, $\widetilde{\beta}_{G}=C_{2}d_{2}^{-\mu_{2}}$, where we have $C_{1}=28$ dB and $C_{2}=26$ dB. Moreover, $\mu_{1}=3.67$ and $\mu_{2}=2.2$ are path-loss components. $d_{1}$ and $d_{2}$ are the distances of BS-users and BS-RIS with $d_{1}=50$ m and $d_{2}=8$ m. Also, we assume that $\eta_{k}$ and $\widetilde{\beta}_{a}$ have the same value as $\widetilde{\beta}_{d,k}$. Excluding the comparison of varying values for these two parameters, it is specified that the system employs $N_{s}=64$ subcarriers with a subcarrier spacing of $\Delta f=240$ kHz (5G NR standard). The duration of an OFDM symbol is $T_{s}=17.68\,\mu$s. The noise powers are given by $\sigma^{2}=\sigma_{v}^{2}=-40$ dB and $\sigma_{r}^{2}=-58$ dB. The maximum value of AAC is $\alpha_{max}=4$. The dimensions for each element within ASTARS are set as $d_{H}=d_{V}=\lambda/4$. We apply the method described in Secs. III and IV to optimize the transmit beamforming and phase coefficients of the ASTARS. By using the results from this optimization, we depict the radar SNR and perform estimation of range and velocity. We assume that the unit of range is m (meters) and the unit of velocity is km/h (kilometers per hour).
VI-A Radar SNR versus number of RIS elements or DFRC BS antennas
@@
In Figs. $2$ and $3$, we have plotted the radar SNR against the number of ASTARS elements and DFRC BS antennas by solving the QSDP problem outlined in Sec. III and Sec. IV. It is clear that when we increase the number of ASTARS panels, the radar SNR increases. This improvement can be attributed to the enhanced reflective capabilities of a larger ASTARS panel, which effectively amplifies the signal reflected back to the radar receiver. With more panels, the ASTARS can reflect more beams towards the target, thereby increasing the signal power at the receiver. Additionally, a larger ASTARS array can better mitigate the interference in our system, further boosting the received signal strength and enhancing the radar SNR. For the number of DFRC BS antennas, we can also find that the radar SNR increases when the number of BS antennas increases. This is due to the beamforming gain achieved with a larger antenna array at the DFRC BS. As the number of antennas increases, the BS can transmit more beams to the ASTARS, which results in a more concentrated energy transmission. Additionally, a larger antenna array provides better spatial resolution and diversity, which further enhances the capability of the system to distinguish between different targets and interference, leading to improved radar detection.
VI-B Radar SNR for ASTARS and STAR-RIS
@@
VI-C Range and velocity MSE for random and optimal ASTARS
Although our optimization approach in Sec. III and Sec. IV primarily aims to maximize the radar SNR, it also enhances range and velocity estimation performance in the sensing domain. It is evident that our optimized ASTARS configuration results in reduced MSE for both range and velocity estimations in Figs. $4$ and $5$. Moreover, as we increase the number of ASTARS elements, the MSEs for both estimations consistently decrease for our optimized case. Conversely, in the random case, MSEs initially drop but subsequently rise, indicating the optimized system capacity to stabilize estimations. This stability is because the optimized ASTARS can concentrate the signal energy more effectively towards the target, thereby improving the measurement accuracy of the signals. By optimizing for radar SNR, we inherently enhance the system ability to distinguish between the signal and noise, which is important for accurate range and velocity estimations. Essentially, a higher radar SNR implies a signal reflection from the target, which directly contributes to more precise sensing outcomes.
VI-D Range and velocity estimations performance analysis
The MSE for range and velocity estimations versus number of ASTARS elements has been shown in Figs. $6$ and $7$. We can find that when we increase the elements number, the MSE decreases at the same time. This is because a more large ASTARS array allows for a more precise manipulation of the phase and amplitude of the reflected signals, thereby improving the accuracy of the phase measurements used in range and velocity calculations. Hence, the estimation performance can also been improved.
In Figs. $6$ and $7$, we can also find that the MSEs for range and velocity estimations increase when their true values increase. Larger distances or higher velocities introduce greater phase shifts across the signal path. Also, as the target moves farther away or at a faster velocity, the signal undergoes more significant changes which can amplify the estimation errors and system imperfections.
VI-E The effect of different number of OFDM subcarriers in the velocity estimation

### v3

@@ I-B Paper outline
For the rest of this paper, Section II introduces the structure of the ASTARS-aided ISAC system and the expressions of radar and communication signals. In Section III, we introduce the optimization problem of our system and the design of transmit beamforming, as well as the optimization of the ASTARS phases and amplitudes. We simplify the problem to enable the use of CVX for carrying out the optimization. In Section IV, we introduce the procedures for range and velocity estimations by using our optimized transmit beamforming and ASTARS phases and amplitudes. Section V includes simulations, and the Conclusion is presented in Section VI.
@@ II-A ASTARS model
We consider a downlink ASTARS-aided ISAC system designed for simultaneous radar sensing and communication as Fig. $1$  [^1] Placing the target and users on opposite sides of the STAR-RIS allows simultaneous optimization of radar sensing and communication performance, leveraging the reflective and transmissive properties of the surface for enhanced system efficiency, see [22, 24].. The DFRC BS transmits a signal to the ASTARS with $Q$ elements, which then directs the beam towards a target in the sensing area. Upon receiving the beam, the target reflects it back along its path, which is then captured by the ASTARS and redirected to the DFRC BS for processing. The DFRC BS employs uniform linear arrays (ULAs) for both sensing and communication with $M$ antennas. These arrays are utilized to communicate with $K$ users, each equipped with a single antenna. Notably, these users are situated on the opposite side of the ASTARS, distinct from the sensing target location. We assume that $\mathcal{K}_{t}=\{1,\ldots,K\}$ users are located on the communication region $(t)$, i.e., behind the surface. We assume that all surface elements serve simultaneously all users in $t$ region and the target in $r$ region. The amplitude and phase parameters for the ASTARS elements are defined as follows: The amplitude for $q$th element $\beta_{q}^{i}$ lies within the range $[0,1]$ and the phase $\phi_{q}^{i}$ within $[0,2\pi)$, where the index $i$ denotes either communication ($t$) or sensing ($r$). The entire phase matrix $\boldsymbol{\Psi}_{i}$ is constructed as a diagonal matrix with elements $\alpha_{q}\beta_{q}^{i}e^{j\phi_{q}^{i}}$, resulting in $\boldsymbol{\Psi}_{i}=\mathrm{diag}(\alpha_{1}\beta_{1}^{i}e^{j\phi_{1}^{i}},\ldots,\alpha_{Q}\beta_{Q}^{i}e^{j\phi_{Q}^{i}})\in\mathbb{C}^{Q\times Q}$. The vector $\boldsymbol{\phi}_{i}$ contains the phase elements of each ASTARS element, expressed as $\boldsymbol{\phi}_{i}=[\alpha_{1}\beta_{1}^{i}e^{j\phi_{1}^{i}},\ldots,\alpha_{Q}\beta_{Q}^{i}e^{j\phi_{Q}^{i}}]^{T}$. It is subject to the conditions that for each element $q$, the sum of the squares of the amplitudes for communication and sensing must be equal to one, i.e., ($(\beta_{q}^{t})^{2}+(\beta_{q}^{r})^{2}=1$), the active amplifying coefficient (AAC) $\alpha_{q}$ should not exceed a maximum value $\alpha_{max}$, and the phase $\phi_{q}^{i}$ must fall within the range $[0,2\pi)$ for all $q$. To acquire the OFDM symbol, the transmitted block signals undergo an inverse discrete Fourier transform (IDFT). Following the IDFT, the output is cyclically shifted to mitigate intersymbol interference (ISI) and to incorporate a guard interval. Once the signals traverse multipath channels, it becomes essential to eliminate the guard interval and execute the discrete Fourier transform (DFT) [28], [31]. The assumption here is that the cyclic prefix (CP) length is sufficiently large to disregard any timing and frequency discrepancies.
@@ II-C Signal at the radar receiver
For the channels in the radar sensing area, we consider that the channel between DFRC BS and the ASTARS is $\mathbf{G}(n)\in\mathbb{C}^{Q\times M}$, which is based on Rician fading with the LoS part and the Rician factor is $R$. The expression of $\mathbf{G}(n)$ can be written as
$\mathbf{G}(n)=\sqrt{\frac{R}{R+1}}\mathbf{G}_{\textup{LoS}}(n)+\sqrt{\frac{1}{R+1}}\mathbf{G}_{\textup{nLoS}}(n),$ (3)
where $\mathbf{G}_{\textup{LoS}}$ and $\mathbf{G}_{\textup{nLoS}}$ are LoS and nLoS parts of the channel $\mathbf{G}$. The LoS $\mathbf{G}_{\textup{LoS}}(n)$ can be written by using steering vectors as $\mathbf{G}_{\textup{LoS}}(n)=\mathbf{a}_{q}(\theta_{0})\mathbf{a}_{m}^{H}(\theta_{0})\in\mathbb{C}^{Q\times M}$ with $\mathbf{a}_{q}(\theta_{0})=[1,e^{-j\pi\textup{sin}(\theta_{0})},\ldots,e^{-j\pi(Q-1)\textup{sin}(\theta_{0})}]^{T}$ and $\mathbf{a}_{m}(\theta_{0})=[1,e^{-j\pi\textup{sin}(\theta_{0})},\ldots,e^{-j\pi(M-1)\textup{sin}(\theta_{0})}]^{T}$. The angle $\theta_{0}$ is from $-\frac{\pi}{2}$ to $\frac{\pi}{2}$. The channel between the ASTARS and the target is $\mathbf{h}_{a,T}(n)\in\mathbb{C}^{Q\times 1}$. To simpify the complexity, the direct link from the DFRC BS to the target is blocked  [^2] Isolating the DFRC BS from the direct link to the target enhances system security and reduces interference, allowing for more accurate radar sensing and communication through the ASTARS, see [21].. Hence, the collected echo signal at the DFRC BS from the target can be written as
$\mathbf{y}_{r}(n)=\mathbf{G}^{H}(n)[\boldsymbol{\Psi}_{r}\mathbf{h}_{a,T}(n)\mathbf{h}_{a,T}^{H}(n)\boldsymbol{\Psi}_{r}\mathbf{G}(n)(\mathbf{x}(n)+\mathbf{v}_{1}(n))$ (4)
@@ II-D Communication signal at the users
$=(\mathbf{t}_{k}^{H}(n)\boldsymbol{\Psi}_{t}^{H}(n)\mathbf{G}(n)+\mathbf{h}_{d,k}^{H}(n))\mathbf{x}(n)+\mathbf{t}_{k}^{H}(n)\boldsymbol{\Psi}_{t}^{H}\mathbf{v}_{1}(n)$
@@
$\mathrm{SINR}_{k}(n)=\frac{|\mathbf{u}_{k}^{H}(n)\mathbf{w}_{c,k}|_{2}^{2}}{|\mathbf{u}_{k}^{H}(n)\mathbf{W}_{r}|_{2}^{2}+\sum_{i\neq k}\mathbf{u}_{k}^{H}(n)\mathbf{R}_{i}\mathbf{u}_{k}(n)+\sigma_{v}^{2}\mathbf{t}_{k}^{H}(n)\boldsymbol{\Psi}_{t}\boldsymbol{\Psi}_{t}^{H}(n)\mathbf{t}_{k}(n)+\sigma^{2}}$ (9)
where $\mathbf{h}_{d,k}(n)\in\mathbb{C}^{M\times 1}$ is the direct link from DFRC BS to users in the $n$th subcarrier; $\mathbf{t}_{k}(n)\in\mathbb{C}^{Q\times 1}$ is the channel between the STAR-RIS and user $k$. $n_{k}(n)\sim\mathcal{CN}(0,\sigma^{2})$ is the AWGN. The SINR for the $k$th user can be given as (9), where $\mathbf{u}_{k}(n)=\mathbf{h}_{d,k}(n)+\mathbf{G}^{H}(n)\boldsymbol{\Psi}_{t}\mathbf{t}_{k}(n)$ is the overall channel between users and the DFRC BS. Also, channels $\mathbf{h}_{d,k}$ and $\mathbf{t}_{k}$ are based on Rayleigh fading, with large-scale fading coefficients $\widetilde{\beta}_{d,k}$, $\widetilde{\beta}_{G}$, and $\eta_{k}$.
@@ III-A Problem formulation
$\quad s.t.\quad\,\,\quad\quad{\color[rgb]{0,0,0}\mathrm{SINR}_{k}(n)\geq\xi,\forall k}$ (10a)
$\quad\quad\quad\quad\quad\quad{\color[rgb]{0,0,0}\mathrm{tr}(\mathbf{W}\mathbf{W}^{H})\leq P_{BS}}$ (10b)
@@ III-B Transmit beamforming $\mathbf{W}$ optimization
$\;\quad{\color[rgb]{0,0,0}\mathrm{tr}(\mathbf{W}\mathbf{W}^{H})\leq P_{BS}},$
@@
III-C Phase coefficients optimization
@@
IV Range and velocity estimation
@@
IV-A Estimations for range and velocity
@@
IV-B MSE for range and velocity estimations
@@
IV-C Complexity of Algorithm $1$
The complexity of Algorithm $1$ primarily depends on two main optimization steps: the transmit beamforming optimization and the ASTARS phase coefficients optimization. The transmit beamforming optimization involves solving a QSDP for the transmit beamforming matrix $\mathbf{W}$, which has a complexity of $\mathcal{O}((M^{2}+MK)^{3.5})$, where the factor $3.5$ reflects the worst-case polynomial time complexity for such problems, taking into account the matrix dimensions of $M$ and $K$. The ASTARS phase coefficients optimization involves converting the optimization problem to a convex form and solving for the phase coefficients $\boldsymbol{\Psi}_{r}$ and $\boldsymbol{\Psi}_{t}$, with a complexity of $\mathcal{O}(Q^{3})$, where the factor $3$ is derived from the cubic time complexity for solving such problems in terms of the number of ASTARS elements $Q$. The total complexity of Algorithm 1 is $\mathcal{O}(((M^{2}+MK)^{3.5}+Q^{3}))$. This reflects the computational intensity of the optimization steps involved.
V Simulations
@@
In this section, we provide simulation results for our proposed ASTARS-aided ISAC systems with OFDM. The DFRC BS is equipped with $M=32$ antennas, which serves $K=4$ users located on the communication area. The ASTARS comprises an equal number of elements in both dimensions, for example, forming a $10\times 10$ array when $Q=100$ elements. In our system, we assume a carrier frequency $3$ GHz with wavelength $\lambda=0.1$ m. We assume the large-scale fading coefficients $\widetilde{\beta}_{d,k}=C_{1}d_{1}^{-\mu_{1}}$, $\widetilde{\beta}_{G}=C_{2}d_{2}^{-\mu_{2}}$, where we have $C_{1}=28$ dB and $C_{2}=26$ dB. To ensure the path loss coefficients are appropriately scaled and less than 1, we convert these values to a linear scale as $C_{1}^{\prime}=10^{-2.8}\approx 1.58\times 10^{-3}$ and $C_{2}^{\prime}=10^{-2.6}\approx 2.51\times 10^{-3}$, maintaining the same path loss characteristics. Moreover, $\mu_{1}=3.67$ and $\mu_{2}=2.2$ are path-loss components. $d_{1}$ and $d_{2}$ are the distances of BS-users and BS-RIS with $d_{1}=50$ m and $d_{2}=8$ m. Also, we assume that $\eta_{k}$ and $\widetilde{\beta}_{a}$ have the same value as $\widetilde{\beta}_{d,k}$. Excluding the comparison of varying values for these two parameters, it is specified that the system employs $N_{s}=64$ subcarriers with a subcarrier spacing of $\Delta f=120$ kHz (5G NR standard). The duration of an OFDM symbol is $T_{s}=17.68\,\mu$s. The noise powers are given by $\sigma^{2}=\sigma_{v}^{2}=-40$ dB and $\sigma_{r}^{2}=-58$ dB. The maximum value of AAC is $\alpha_{max}=4$. The dimensions for each element within ASTARS are set as $d_{H}=d_{V}=\lambda/4$. The channel $\mathbf{G}$ is based on Rician fading with Rician factor $R=10$ dB. We apply the method described in Sec. III to optimize the transmit beamforming and phase coefficients of the ASTARS. By using the results from this optimization, we depict the radar SNR and perform estimation of range and velocity. We assume that the unit of range is m (meters) and the unit of velocity is km/h (kilometers per hour).
V-A Radar SNR versus number of RIS elements or DFRC BS antennas
@@
In Figs. $2$ and $3$, we have plotted the radar SNR against the number of ASTARS elements and DFRC BS antennas by solving the QSDP problem outlined in Sec. III. It is clear that when we increase the number of ASTARS panels, the radar SNR increases. This improvement can be attributed to the enhanced reflective capabilities of a larger ASTARS panel, which effectively amplifies the signal reflected back to the radar receiver. With more panels, the ASTARS can reflect more beams towards the target, thereby increasing the signal power at the receiver. Additionally, a larger ASTARS array can better mitigate the interference in our system, further boosting the received signal strength and enhancing the radar SNR. For the number of DFRC BS antennas, we can also find that the radar SNR increases when the number of BS antennas increases. This is due to the beamforming gain achieved with a larger antenna array at the DFRC BS. As the number of antennas increases, the BS can transmit more beams to the ASTARS, which results in a more concentrated energy transmission. Additionally, a larger antenna array provides better spatial resolution and diversity, which further enhances the capability of the system to distinguish between different targets and interference, leading to improved radar detection.
V-B Radar SNR for ASTARS and STAR-RIS
@@
V-C Range and velocity MSE for random and optimal ASTARS
Although our optimization approach in Sec. III primarily aims to maximize the radar SNR, it also enhances range and velocity estimation performance in the sensing domain. It is evident that our optimized ASTARS configuration results in reduced MSE for both range and velocity estimations in Figs. $4$ and $5$. Moreover, as we increase the number of ASTARS elements, the MSEs for both estimations consistently decrease for our optimized case. Conversely, in the random case (RIS phase is random from $-\frac{\pi}{2}$ to $\frac{\pi}{2}$), MSEs initially drop but subsequently rise, indicating the optimized system capacity to stabilize estimations. This stability is because the optimized ASTARS can concentrate the signal energy more effectively towards the target, thereby improving the measurement accuracy of the signals. By optimizing for radar SNR, we inherently enhance the system ability to distinguish between the signal and noise, which is important for accurate range and velocity estimations. Essentially, a higher radar SNR implies a signal reflection from the target, which directly contributes to more precise sensing outcomes.
V-D Range and velocity estimations performance analysis
The MSE for range and velocity estimations versus number of ASTARS elements has been shown in Figs. $6$ and $7$. We can find that when we increase the elements number, the MSE decreases at the same time. This is because a more large ASTARS array allows for a more precise manipulation of the phase and amplitude of the reflected signals, thereby improving the accuracy of the phase measurements used in range and velocity calculations. Hence, the estimation performance can also been improved. In Figs. $6$ and $7$, we can also find that the MSEs for range and velocity estimations increase when their true values increase. Larger distances or higher velocities introduce greater phase shifts across the signal path. Also, as the target moves farther away or at a faster velocity, the signal undergoes more significant changes which can amplify the estimation errors and system imperfections.
V-E The effect of different number of OFDM subcarriers in the velocity estimation